OpenStudy (natasha18):

Given the sequence 2, 6, 18, 54, ..., which expression shown would give the fifteenth term? 2^15 2·3^15 2·3^14

OpenStudy (natasha18):

I think it's the second one..

OpenStudy (natasha18):

Right?

OpenStudy (decarr432):

we can rule a out easily now work out b and let me know when you finish

OpenStudy (natasha18):

28,697,814 is what I got

OpenStudy (albert0898):

Okay, you need to know one thing to answer this question. 1) Geometric Sequence The Geometric Sequence Formula is \[a_n = a _{1} * r ^{n-1}\] an = number we're looking for a1 = 2 r = ration = 3 n = number of terms = 15 \[a_n = 2 * 3 ^{14}\]

OpenStudy (boldjon):

2,6,18,54 This is a geometric sequence since there is a common ratio between each term. In this case, multiplying the previous term in the sequence by 3 gives the next term. In other words, an=r*an-1. Geometric Sequence: r=3 Remove the parentheses around the expression 15. a15=2*3^(15-1) Subtract 1 from 15 to get 14. a15=2*3^(14) Simplify the result. a15=2*(4782969) Multiply 2 by each term inside the parentheses. a15=9565938

OpenStudy (boldjon):

so D

OpenStudy (solomonzelman):

Generally, as regards to "geometric sequences" with "common ratio" r, and the first term \(a_1\) we can write the equation for "any nth term" which is denoted as, \(a_n\). \(\color{#000000 }{ \displaystyle a_n=a_1\times r^{n-1} }\) Or, you know that any "(n-1)st term" multiplied times the common ratio "r", gives you the "nth term" (a term after (n-1)st term). Or, equation-wise, \(\color{#000000 }{ \displaystyle a_n=a_{n-1}\times r }\) For example, when this last formula is applied, \(\color{#000000 }{ \displaystyle a_2=a_1\times r }\) and that can be rearranged to \(\color{#000000 }{ \displaystyle a_{2}/ r=a_2\cdot r^{-1}=a_1 }\) So you can write, \(\color{#000000 }{ \displaystyle a_n=a_2\cdot r^{-1}\times r^{n-1}=a_2 \cdot r^{n-2} }\) In fact, if you think about it, \(\color{#000000 }{ \displaystyle a_n=a_k\cdot r^{-1}\times r^{n-k} }\) for any integer k. So in summary, \(\color{#000000 }{ \displaystyle a_n=a_{n-1}\times r }\) \(\color{#000000 }{ \displaystyle a_n=a_1\times r^{n-1} }\) \(\color{#000000 }{ \displaystyle a_n=a_k\cdot r^{-1}\times r^{n-k} }\)

OpenStudy (solomonzelman):

Oh, when I said "in fact ..." the equation there should be just \(\color{#000000 }{ \displaystyle a_n=a_k\cdot r^{n-k} }\)

OpenStudy (solomonzelman):

And so is the last equation in the post, just, \(\color{#000000 }{ \displaystyle a_n=a_k\cdot r^{n-k} }\)

OpenStudy (natasha18):

Thanks everyone for the help! I get it now :)

OpenStudy (dangalique):

I think that its the second one.